OpenStudy (anonymous):

Find the solutions of the following system of equations. y = x² + 3x – 4 y = 2x + 2

OpenStudy (anonymous):

If you factor the top equation you find it has only 2 roots. 1 and -3. Only 1 will work in the second one

OpenStudy (anonymous):

y = x² + 3x – 4 y = 2x + 2 2x + 2 = x² + 3x – 4 x² + x -6 =0 (x+3)(x-2)=0 x=-3 or 2 for x=-3, y=2(-3)+2 = -4 for x=2, y=2(2)+2 = 6

OpenStudy (anonymous):

whoops :S

OpenStudy (anonymous):

@mcnalljj: That is not one of my answer choices

OpenStudy (anonymous):

:P that's cause it's wrong...ignore me follow CoCoTsoi's method